OpenStudy (anonymous):

Rewrite the equation of a line in point-slope form to slope intercept form: y-2=2/3(6x+9)

OpenStudy (anonymous):

You have this: y-2=2/3(6x+9) You want to transform that into this: y=mx+b What do you think the first step should be?

OpenStudy (anonymous):

Simplify so +2?

OpenStudy (anonymous):

Not quite yet. First you want to distribute 2/3 into (6x+9) Do you remember how to do that?

OpenStudy (anonymous):

yes so 4x+6?

OpenStudy (anonymous):

Yep! y-2=4x+6 Looks much better now, right? NOW we can add two to both sides. How would the equation look after we do that?

OpenStudy (anonymous):

It would be y=4x+8

OpenStudy (anonymous):

Correct! And since the equation is now in slope intercept form, that's your answer!

OpenStudy (anonymous):

Thank you! but theres a few more, if you could help that would be greaaaat!

OpenStudy (anonymous):

If I know how to do it, I'll help!

OpenStudy (anonymous):

Alright. So you know what a y-intercept is, correct?

OpenStudy (anonymous):

Somewhat

OpenStudy (anonymous):

Well all you need to know for now is that the y-intercept is where the line crosses the y-axis. Looking at the graph, can you tell what the y-intercept is?

OpenStudy (anonymous):

0,2

OpenStudy (anonymous):

Yes. But we can just call it 2. So let's plug that into the slope intercept formula. y = mx + 2 Next, we want to find m. What would we look at to find that?

OpenStudy (anonymous):

The x intercept?

OpenStudy (anonymous):

Not quite.. We would look at the slope of the line. Also known as rise/run. What is the rise/run of this graph?

OpenStudy (anonymous):

I honestly dont know :s

OpenStudy (anonymous):

Okay.. The rise of a line is how much distance up/downward the line moves to get to the next point. Run is the distance left/right. Like this.

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

yes but there arent any points

OpenStudy (anonymous):

or is it 10,3?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

Sure there are. Points are formed whenever a line crosses defined coordinates. The next point up would be at (1, 5).

OpenStudy (anonymous):

....? Im losyt

OpenStudy (anonymous):

I'm sorry. I guess I'm not explaining this clearly enough. A good way to spot where the next "point" on a line is, is to look wherever the line passes through the center of a "+" on the grid. Is that understandable, or is it still confusing?

OpenStudy (anonymous):

Yes, I understand, what would I need to do next?

OpenStudy (anonymous):

We still need to find the rise and the run. As you can see on the graph, the distance between (0, 2), and (1, 5) is 3 up, 1 across. Making the slope 3/1, which is 3. So where would that go in the formula y = mx + 2 ?

OpenStudy (anonymous):

m=3

OpenStudy (anonymous):

Yep! And that's the equation of the line in slope intercept form! y = 3x + 2

OpenStudy (anonymous):

thankyou, I need help with just 3 more

OpenStudy (anonymous):

Write the equation of a line in slope intercept form that is parallel to the line y=1/3x+5 and passes through (-9, 5).

OpenStudy (anonymous):

Okay. So here are both of our equations at this point. y = (1/3)x+5 y = mx + b

OpenStudy (anonymous):

So as we discussed earlier, if two lines are parallel, then they share the same slope. So plug the slope into the second equation.

OpenStudy (anonymous):

y=5/(1/3)?

OpenStudy (anonymous):

Hmm?

OpenStudy (anonymous):

...

OpenStudy (anonymous):

I'm not sure what you did there. The equation that we are trying to find has to have the same slope as the equation we are given. (1/3) Plug (1/3) into y = mx+b

OpenStudy (anonymous):

1/3=b?

OpenStudy (anonymous):

or m?

OpenStudy (anonymous):

m is always the slope. (1/3) = m So we have y = (1/3)x + b We need to find b. But we are already given the x and y values (-9, 5). Plug those into the equation we already have. y=(1/3)x+b

OpenStudy (anonymous):

y=(1/3)x+-4?

OpenStudy (anonymous):

No. Don't add the two values together, just plug them into the equation.. 5 = (1/3)(-9) + b Now solve the equation for b.

OpenStudy (anonymous):

b=5

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

I'm guessing that you aren't understanding the things that I am explaining?

OpenStudy (anonymous):

If not, it's fine. I have been doing this for quite a bit, so this may be a sign that I need to take a break.

OpenStudy (anonymous):

But here is the rest of the problem. You multiply (1/3) and (-9) to get -3, then you add that to both sides to get b = 8. We then get the equation y =(1/3)x + 8

OpenStudy (anonymous):

And that's your answer.